find three consecutive integers such that the difference of the squares of the least and the greatest is 96
Answer by stanbon(75887)

find three consecutive integers such that the difference of the squares of the least and the greatest is 96
------
1st: x-1
2nd: x
3rd: x+1
-------
Equation:
(x-1)^2-(x+1)^2 = 96
======
x^2 -2x + 1 - (x^2 +2x + 1) = 96
-4x = 96
x = -24
---
1st: -25
2nd: -24
3rd: -23
-------
Cheers,
Stan H.
================